# Powered Descent Trajectory Optimization of Chandrayaan-3 using Radau Collocation and Controllable Sets

### v1

@@ I INTRODUCTION
Trajectory design involves optimizing a performance metric while satisfying state and control constraints. As a result, it is often formulated as an optimal control problem, also referred to as trajectory optimization. With the exception of simple problems such as infinite-horizon LQR, optimal control problems must be solved numerically. The need to solve optimal control problems numerically has led to a wide range of numerical approaches. These numerical approaches are broadly classified into two categories: indirect methods, which rely on solving the necessary conditions of optimality, and direct methods, which transcribe the optimal control problem into a nonlinear program. A survey of different methods for optimal control is provided in [1]. Collocation methods are among the most widely used direct approaches for trajectory optimization. Their popularity has grown with the rapid advancement of efficient solvers capable of handling large-scale nonlinear program[2, 3, 4]. The two most common forms of collocation are local collocation and global collocation. Local collocation employs Runge-kutta methods[5] and orthogonal collocation to discretize the optimal control problem[6]. The key difference between an orthogonal collocation and Runge-kutta method is the manner in which the collocation points are chosen. In orthogonal collocation, the collocation points are the roots of an orthogonal polynomial associated with a quadrature rule to approximate definite integrals. In recent years, pseudospectral methods have gained popularity. They are a global form of orthogonal collocation, where the state is approximated by a global polynomial, and collocation is applied at selected points. Unlike local collocation, which fixes the polynomial degree and varies the number of meshes, pseudospectral methods fix the number of meshes and vary the polynomial degree [7, 8].
@@ II-A Multiphase Optimal Control Problem
Here, $r_{s},l_{s}\in\{1,\ldots,S\}$ are the right and left phases, respectively, of the linkage pairs, with $r_{s}\neq l_{s}$ (implying that a phase cannot be linked to itself) and $S$ is the number of pairs of phases to be linked. While [1] considers more general linkage constraints, in this work we restrict them to sequential linkages, which implies which implies $r_{s}>l_{s}$.
@@ III-A Forward Pass - Trajectory Optimization
Finally, the waypoints obtained from trajectory optimization are implemented through a suitable guidance algorithm in real time. Perhaps, the most widely cited approach for solving the fuel-optimal powered descent guidance problem is lossless convexification [9], which transforms the original non-convex fuel optimal control problem into a convex optimization problem that can be solved onboard in real time. Although such methods explicitly handle constraints, the computational burden on low-speed processors limits their real-time applicability. Consequently, onboard guidance laws typically rely on analytical formulation under simplifying assumptions, often neglecting state and actuation constraints, and thus yield suboptimal guidance law. Therefore, it is essential to evaluate the robustness of the suboptimal guidance law when applied to the optimized waypoints. This necessitates a backward pass in the trajectory design process, wherein the waypoints are further refined by taking into account the specific characteristics and limitations of the onboard guidance law.
@@ III-B Backward Pass - Controllability based Refinement
Fig 4 shows the dataset in this reduced state space. The point labeled O represents the optimized fine braking waypoint obtained from the forward pass. Since dispersion in optimized state is assumed to follow a Gaussian process, the controllability boundary is parameterized using a convex conic function using maximum margin classifier [10] as
@@
$\max_{S,v,w}\;\mathcal{J}(S,v,w)=-\lambda M(S/v,H/w)+(1-\lambda)\Delta m^{*}(S,H,v,w)$ (38)
@@ III-C Terminal Descent Trajectory Design
The terminal descent phase starts with 12 sec hovering where two engines are shut down to ensure guidance required thrust is within engines operational range. lander navigation is switched from inertial reference to terrain relative frame centered below the lander on the local terrain. At the end of the first hovering, guidance uses the latest navigation estimates to plan the vertical descent to 150 m. Vertical descent phase ends at 150 m altitude with a planned second hovering phase for hazard detection. The second hovering phase lasts for a maximum of 22 seconds. During this phase, the hazard camera is operated to provide hazard-free coordinates with respect to current location. Once a hazard free landing point is identified, hovering is terminated and retargeting to safe location is initiated with latest states estimated from navigation. Retargeting is done to an altitude of 60 m above hazard free landing location ensuring collision free retargeting. After retargeting, the lander descends to 10 meters, aiming for a target vertical velocity of -1 m/s, followed by a constant velocity of -1 m/s descent until touchdown sensor detects surface contact with footpad. Fig. 6 shows the various waypoints obtained after the trajectory design for the final flight. More details on the landing profile and powered descent trajectory for Chandrayaan-3 can be found in [11].

### v2

@@ I INTRODUCTION
Trajectory design involves optimizing a performance metric while satisfying state and control constraints. As a result, it is often formulated as an optimal control problem, also referred to as trajectory optimization. With the exception of simple problems such as infinite-horizon LQR, optimal control problems must be solved numerically. The need to solve optimal control problems numerically has led to a wide range of numerical approaches. These numerical approaches are broadly classified into two categories: indirect methods, which rely on solving the necessary conditions of optimality, and direct methods, which transcribe the optimal control problem into a nonlinear program. A survey of different methods for optimal control is provided in [rao2009survey]. Collocation methods are among the most widely used direct approaches for trajectory optimization. Their popularity has grown with the rapid advancement of efficient solvers capable of handling large-scale nonlinear program[betts1992application, byrd1999interior, vanderbei1999interior]. The two most common forms of collocation are local collocation and global collocation. Local collocation employs Runge-kutta methods[dontchev2000second] and orthogonal collocation to discretize the optimal control problem[reddien1979collocation]. The key difference between an orthogonal collocation and Runge-kutta method is the manner in which the collocation points are chosen. In orthogonal collocation, the collocation points are the roots of an orthogonal polynomial associated with a quadrature rule to approximate definite integrals. In recent years, pseudospectral methods have gained popularity. They are a global form of orthogonal collocation, where the state is approximated by a global polynomial, and collocation is applied at selected points. Unlike local collocation, which fixes the polynomial degree and varies the number of meshes, pseudospectral methods fix the number of meshes and vary the polynomial degree [garg2010unified, garg2011direct].
@@ II-A Multiphase Optimal Control Problem
Here, $r_{s},l_{s}\in\{1,\ldots,S\}$ are the right and left phases, respectively, of the linkage pairs, with $r_{s}\neq l_{s}$ (implying that a phase cannot be linked to itself) and $S$ is the number of pairs of phases to be linked. While [rao2009survey] considers more general linkage constraints, in this work we restrict them to sequential linkages, which implies which implies $r_{s}>l_{s}$.
@@ III-A Forward Pass - Trajectory Optimization
Finally, the waypoints obtained from trajectory optimization are implemented through a suitable guidance algorithm in real time. Perhaps, the most widely cited approach for solving the fuel-optimal powered descent guidance problem is lossless convexification [malyuta2022convex], which transforms the original non-convex fuel optimal control problem into a convex optimization problem that can be solved onboard in real time. Although such methods explicitly handle constraints, the computational burden on low-speed processors limits their real-time applicability. Consequently, onboard guidance laws typically rely on analytical formulation under simplifying assumptions, often neglecting state and actuation constraints, and thus yield suboptimal guidance law. Therefore, it is essential to evaluate the robustness of the suboptimal guidance law when applied to the optimized waypoints. This necessitates a backward pass in the trajectory design process, wherein the waypoints are further refined by taking into account the specific characteristics and limitations of the onboard guidance law.
@@ III-B Backward Pass - Controllability based Refinement
Fig 4 shows the dataset in this reduced state space. The point labeled O represents the optimized fine braking waypoint obtained from the forward pass. Since dispersion in optimized state is assumed to follow a Gaussian process, the controllability boundary is parameterized using a convex conic function using maximum margin classifier [chakrabarti2024convex] as
@@
$\min_{S,v,w}\;\mathcal{J}(S,v,w)=-\lambda\mathcal{M}(S/v,H/w)+(1-\lambda)\Delta m^{*}(S,H,v,w)$ (38)
@@ III-C Terminal Descent Trajectory Design
The terminal descent phase starts with 12 sec hovering where two engines are shut down to ensure guidance required thrust is within engines operational range. lander navigation is switched from inertial reference to terrain relative frame centered below the lander on the local terrain. At the end of the first hovering, guidance uses the latest navigation estimates to plan the vertical descent to 150 m. Vertical descent phase ends at 150 m altitude with a planned second hovering phase for hazard detection. The second hovering phase lasts for a maximum of 22 seconds. During this phase, the hazard camera is operated to provide hazard-free coordinates with respect to current location. Once a hazard free landing point is identified, hovering is terminated and retargeting to safe location is initiated with latest states estimated from navigation. Retargeting is done to an altitude of 60 m above hazard free landing location ensuring collision free retargeting. After retargeting, the lander descends to 10 meters, aiming for a target vertical velocity of -1 m/s, followed by a constant velocity of -1 m/s descent until touchdown sensor detects surface contact with footpad. Fig. 6 shows the various waypoints obtained after the trajectory design for the final flight. More details on the landing profile and powered descent trajectory for Chandrayaan-3 can be found in [rallapalli2024landing].